If y(x, t) = (5.2 mm) sin(kx + (520 rad/s)t + ϕ) describes a wave traveling along a string, how much time does any given point on the string take to move between displacements y = -2.0 mm and y = +2.0 mm?

Can someone do this, and explain how? I have a quiz on this subject matter soon and I have absolutely no idea how to do this, and I have almost no understanding of what is going on in this problem. Thank you in advance for any help you provide!
Answer #1

when y=-2 mm

-2=5.2Sinleft ( Kx+520t +phi ight )

(금) > Kr +520t +0= Sin 一一0.3948 5.2

when y=2 mm

2=5.2Sinleft ( Kx+520t +phi ight )

-G)-0.3948 (흙 > Kr +520t +0= Sin 5.2

2-1

=>Kx+520t+phi-Kx-520t-phi =0.3948-(-0.3948)

1040t-0.7896

t=7.592*10-4 s (or) 0.7592 ms